OpenStudy (anonymous):

The ratio of a to b is 4 to 5 and the ratio to a to c is 2 to 7 , what is the ratio of c to b ?

OpenStudy (anonymous):

\[\frac{a}{b} = \frac{4}{5}\] \[\frac{a}{c} = \frac{2}{7}\] Just divide them..

OpenStudy (anonymous):

Please show me

OpenStudy (anonymous):

\[\huge \frac{\frac{a}{b}}{\frac{a}{c}} = \frac{\frac{4}{5}}{\frac{2}{7}} \implies \frac{a \times c}{b \times a} = \frac{4 \times 2}{5 \times 7}\]

OpenStudy (anonymous):

No check it once again..

OpenStudy (anonymous):

Thanks I got it 8/35

OpenStudy (anonymous):

On left hand side a and a gets cancel.. So you are left with: \[\frac{c}{b} = \frac{4 \times 2}{5 \times 7}\]

OpenStudy (anonymous):

Yes you are right @Tanzina